OpenStudy (anonymous):

Solve the equation-5y-9=(y-1) answers are. -1/2 , -21/2, -2, -2/4

OpenStudy (anonymous):

none of these are right…. the answer is -4/3

OpenStudy (tkhunny):

-5y - 9 = y - 1 Add 1 -5y - 8 = y Add 5y - 8 = 6y Divide by 2 -4 = 3y Divide by 3 -4/3 = y I dare the "answers are" to refute it.